# Fairness-Regulated Dense Subgraph Discovery

### v1

@@ 4 Selecting the Regularization Parameter $\lambda$
The first step is to determine the start and end points of the regularization paths of problems (1) and (3), which determine the intervals characterizing all possible solutions generated by each formulation. While $\lambda=0$ corresponds to the start of the regularization path for both problems (note that this is tantamount to solving DSG in both cases), determining the end is a non-trivial task in general. This stems from the fact that both problems are combinatorial in nature. In the context of FADSG-I, we wish to determine $\lambda_{\max}$ such that for all $\lambda\geq\lambda_{\max}$, the solution ${\mathbf{x}}^{*}(\lambda)$ is the densest subset of the protected subset $\mathcal{S}_{p}$. Let $\mathcal{S}_{p}^{*}\subseteq\mathcal{S}_{p}$ denote the densest protected subset and $\rho(\mathcal{S}_{p}^{*})$ denote its density. While it is difficult to pin down the exact analytical form of $\lambda_{\max}$, the following result can be derived (see Appendix D.1 for proof).
@@
The above result shows that there indeed exists a finite value of $\lambda$ that corresponds to the end of the regularization path of FADSG-I. However, it is difficult to further simplify the result. Meanwhile, for FADSG-II, the analogous definition of $\lambda_{\max}$ corresponds to the value such that for all $\lambda\geq\lambda_{\max}$, the solution ${\mathbf{x}}^{*}(\lambda)$ is the protected subset $\mathcal{S}_{p}$. We can show that there exists a finite $\lambda$ that demarcates the end of the regularization path for FADSG-II, similarly to FADSG-I (see Appendix D.2). In practice, it is computationally easier to find a suitable $\lambda_{\max}$ (which we know exists owing to Propositions 3 and 5 in Appendix D) via trial and error (see Appendix H, Table 6 for values used). This enables us to bound the interval of the regularization path for a given dataset, which in turn facilitates the selection of the regularization parameter that corresponds to a target fairness level, as explained next.
@@ 4.1 Selection of Target Fairness Level
FADSG-I: For $\lambda=0$, let ${\mathbf{x}}^{*}(0)$ denote the solution of problem (1); i.e., the solution of DSG. We also denote $l_{1}:=r_{1}({\mathbf{x}}^{*}(0))\geq 0$ as the fraction of protected vertices present in the solution of DSG. This corresponds to the level of fairness (as measured by $r_{1}(\cdot)$) that is attainable by the solution of the non-fair DSG problem, and is not guaranteed to be non-zero in general. Meanwhile, we know that ${\mathbf{x}}^{*}(\lambda_{\max})$ corresponds to the densest protected subset for which $u_{1}:=r_{1}({\mathbf{x}}^{*}(\lambda_{\max}))=1$. Let $\alpha\in[l_{1},u_{1}]$ denote the desired fairness level; i.e., we wish to select $\lambda$ such that the solution of (1), ${\mathbf{x}}^{*}(\lambda)$, is the densest subgraph that satisfies $r_{1}({\mathbf{x}}^{*}(\lambda))=\alpha$. To this end, we will exploit the fact that the function $r_{1}({\mathbf{x}}^{*}(\lambda))$ is non-decreasing in $\lambda$. Consequently, if we define the function
@@ 5 Quantifying the Price of Fairness
FADSG-I and II trade off subgraph density in order to promote the inclusion of protected vertices. Consequently, the density of the fair subgraphs obtained via (1) and (3) is no larger than the density of their non-fair counterpart DSG; i.e., a loss in subgraph density is the price paid for ensuring fairness. The loss incurred in attaining a target fairness level $\alpha$, which corresponds to the fraction of protected vertices in the solution, can be quantified using the price of fairness.
Given a graph $\mathcal{G}$, let $l_{1}$ denote the fraction of protected vertices in the solution of DSG (without fairness), and $\alpha\in[l_{1},1]$ the target fairness level. The price of fairness (PoF) for a fixed value of $\alpha$ is
$\textsc{PoF}(\mathcal{G},\alpha):=\frac{\rho_{\mathcal{G}}^{*}-\rho_{\mathcal{G}}^{*}(\alpha)}{\rho_{\mathcal{G}}^{*}}:=1-\frac{\rho_{\mathcal{G}}^{*}(\alpha)}{\rho_{\mathcal{G}}^{*}},$ (6)
@@
where $\rho_{\mathcal{G}}^{*}(\alpha)$ is the density of the densest subgraph in $\mathcal{G}$ containing a fraction $\alpha$ of protected vertices, and $\rho_{\mathcal{G}}^{*}$ is the density of the solution of DSG. Simply stated, $\textsc{PoF}(\mathcal{G},\alpha)$ is the ratio of the loss in subgraph density suffered in meeting the fairness requirement $\alpha$ and the maximum density $\rho_{\mathcal{G}}^{*}$ in the absence of fairness considerations. Note that when $\alpha=l_{1}$, we have $\rho_{\mathcal{G}}^{*}(l_{1})=\rho_{\mathcal{G}}^{*}$, and thus $\textsc{PoF}(\mathcal{G},l_{1})=0$. As $\alpha$ increases, the density $\rho_{\mathcal{G}}^{*}(\alpha)$ decreases $\textsc{PoF}(\mathcal{G},\alpha)$ approaches $1$. For a desired value of $\alpha$, the closer $\textsc{PoF}(\mathcal{G},\alpha)$ is to $1$, the greater is the price (in density) paid for extracting the desired fair dense subgraph. Similar definitions have been considered for analyzing fairness in influence maximization [44] and graph covering [45]. To the best of our knowledge, however, we are the first to study such a notion in the context of detecting fair dense subgraphs. Analyzing the price of fairness is inherently complex due to multiple factors such as the type of graph considered, the underlying protected subgraph, and the desired level of fairness.
@@
We base our study on a difficult example which exemplifies the underlying tension between density and fairness. Consider the “lollipop” graph (denoted as $\mathcal{L}_{n}$). In this construction, the unprotected vertices constitute a clique of size $\sqrt{n}$ ($\mathcal{C}_{\sqrt{n}}$), while the protected vertices form a path graph on $n-\sqrt{n}$ vertices ($\mathcal{P}_{n-\sqrt{n}})$, with one end connected to $\mathcal{C}_{\sqrt{n}}$ (so that the overall graph is connected). We point out that the lollipop graph mimics the properties of real-world graphs in the following manner: (i) the graph $\mathcal{L}_{n}$ has $O(n)$ edges, and is hence sparse, (ii) the majority of vertices ($n-\sqrt{n}$ to be precise) have low degree ($\leq 2$), while the remaining vertices have high degree (at least $\sqrt{n}-1$). Intuitively, we expect the solution of DSG for the lollipop graph to be the clique $\mathcal{C}_{\sqrt{n}}$, as it corresponds to a region of high density. Since the solution of DSG does not contain any protected vertices, from a fairness perspective, it is the worst possible solution. In addition, as the protected subset $\mathcal{P}_{n-\sqrt{n}}$ is minimally connected, and thus has low density, we expect that improving the representation of protected vertices will cost a large decrease in density, which in turn will be reflected by a large price of fairness.
@@
Let us analyse how the PoF is affected by altering the balance between the majority and minority vertices in the desired solution. Consider the case where the desired proportion of unprotected and protected vertices in the extracted subgraph is $1:\gamma$, with $\gamma\in\{1,2,\cdots,n-\sqrt{n}\}$, and the desired balance is $\alpha_{\gamma}:=\gamma/(1+\gamma)$. Note that the case of $\gamma=1$ corresponds to the perfect balance scenario considered in [21, 22]. Then, the PoF can be expressed as follows (see Appendix E for proof).
@@ 6 Algorithmic Approach
An alternative is to consider the iterative convex-optimization algorithms developed in [28, 20, 19]. These methods are centered around solving the dual LP relaxation of DSG proposed in [15]. However, modifying and implementing these algorithms (which are tailored for DSG) for our problems poses a greater challenge relative to Super-Greedy++. First, FADSG-I and II have to be reformulated as a constrained convex optimization problem via an extension of the dual linear programming (LP) relaxation of DSG proposed in [15]. Moreover, since the algorithms of [28, 20, 19] are tailored to exploit the structure of the dual LP relaxation based on DSG, it is not clear to what extent they would need to be modified in order to be applicable for our problems, and whether the requisite changes would still result in low-complexity updates. Another practical drawback of these continuous methods (even when applied to plain DSG) is that a subroutine known as fractional peeling [20] has to be applied on each of the generated iterates if one desires to extract an estimate of the densest supermodular subset (which corresponds to the densest subgraph for DSG) at each iteration. However, performing a single round of fractional peeling incurs a complexity of $O(m+n\log n)$, which has the same complexity order as a single iteration of Super-Greedy++. Consequently, determining the point at which the estimate of the DSS has converged from the continuous iterates can prove to be computationally expensive, unless the fractional peeling subroutine is sparingly used. Even then, this can come at the cost of executing the convex optimization algorithms for a larger number of iterations than required for the sequence of DSS estimates to converge. We point out that this is a non-issue in Super-Greedy++, as its combinatorial nature makes it easy to keep track of the best estimate of the DSS at each iteration. Owing to these reasons, we elect to choose Super-Greedy++ as the algorithmic primitive for FADSG-I and FADSG-II.
@@
Note that the objective function of FADSG-I can be expressed in the form of the DSS problem, with a non-negative, monotone supermodular numerator. Hence, the Super-Greedy++ algorithm can be applied directly to problem (1) to obtain a high-quality approximate solution. For FADSG-II, its objective function can be expressed as well as the ratio $F(\mathcal{S})/|\mathcal{S}|$, where the numerator is supermodular, making problem (3) a special case of DSS. However, since the numerator may not always be non-negative for every $\lambda>0$, Super-Greedy++ does not guarantee an $(1-\epsilon)$-approximation in all cases (see [47] for a counter-example). Nonetheless, Super-Greedy++ can still be applied on problem (3) and our experiments indicate that it can serve as an effective heuristic. For details regarding implementation, refer to Appendix F.
@@ 7 Experimental Evaluation
Experimental Setup: We consider various real-world datasets, with Table 1 summarizing their statistics (for more information, see Appendix G).
@@
Table 1: Summary of dataset statistics: number of vertices ($n$), number of edges ($m$), and size of protected subset ($n_{p}$). The datasets contain multiple graphs and the statistics represent mean values with standard deviations. The full list can be found in Appendix G.
@@
$\bullet$ 2-DFSG [21, 22]. The algorithm starts from the solution of DSG and subsequently introduces vertices from the underrepresented group in an arbitrary fashion, until perfect balance is reached.
$\bullet$ PS, FPS [21, 22]. Paired Sweep (PS) and Fair Paired Sweep (FPS) are spectral algorithms. They categorize the elements based on their group and sort them in a manner reminiscent of spectral clustering.
@@ 7.1 Evaluation of FADSG-I
Comparison with prior art: In order to compare FADSG-I with the chosen baselines, we set the target fairness level $\alpha=0.5$. In this case, our formulation aims to extract a dense subgraph with perfect balance. We point out that we exclude datasets for which the protected vertices in the DSG solution are more than the unprotected, as they already meet and exceed the target $\alpha=0.5$. Figures 2 and 3 illustrate a comparative analysis with prior art. The top sections showcase the price of fairness (PoF) for each algorithm across various datasets, while the bottom sections depict the fraction of protected vertices in the solutions. As the results indicate, our proposed formulation consistently yields the lowest PoF values across a broad spectrum of scenarios. [^1] Note that, in all experiments, in order to compute $\rho^{*}_{\mathcal{G}}$ for the determination of PoF, we employ the exact max flow algorithm [14].
@@
It can be observed from Figures 2 and 3 that for certain datasets, FADSG-I does not find a perfectly balanced subgraph (see Figure 3, Amazon hpc & op). To obtain a better understanding, we plot the regularization path of $r_{1}(\mathcal{S})$ for these datasets in Figure 4. A noteworthy observation is the emergence of “fairness plateaus” where the fairness level remains constant across intervals of $\lambda$. Since FADSG-I is tractable, these figures correspond to the actual Pareto frontier which characterizes the trade-off between fairness and density. This phenomenon suggests that the graph’s inherent structure leads to certain fairness levels that arise naturally. For instances where the data’s underlying structure does not inherently support a perfect division (i.e., Amazon hpc & op), FADSG-I does not enforce it. Figure 5 depicts the outcomes in datasets where there exists an extreme imbalance in the composition of the protected and unprotected subgroups, i.e., we have $n_{p}\ll n$ (see Table 4, Appendix G). In such cases, with the exception of $2$-DFSG, the prior art fails to extract a balanced/fair solution. In fact, we observe an extreme gap to the desired fairness level of $0.5$, as well as a PoF that is, in almost all of the cases, worse than that of our algorithm. This is an important feature of FADSG-I, as it can provide a near optimal solution, when the nature of the graph allows it, while generally paying the least in terms of PoF.

### v2

@@ 4 Selecting the Regularization Parameter $\lambda$
The first step is to determine the start and end points of the regularization paths of problems (2) and (4), which determine the intervals characterizing all possible solutions generated by each formulation. While $\lambda=0$ corresponds to the start of the regularization path for both problems (note that this is tantamount to solving DSG in both cases), determining the end is a non-trivial task in general. This stems from the fact that both problems are combinatorial in nature. In the context of FADSG-I, we wish to determine $\lambda_{\max}$ such that for all $\lambda\geq\lambda_{\max}$, the solution ${\mathbf{x}}^{*}(\lambda)$ is the densest subset of the protected subset $\mathcal{S}_{p}$. Let $\mathcal{S}_{p}^{*}\subseteq\mathcal{S}_{p}$ denote the densest protected subset and $\rho(\mathcal{S}_{p}^{*})$ denote its density. While it is difficult to pin down the exact analytical form of $\lambda_{\max}$, the following result can be derived (see Appendix C.1 for proof).
@@
The above result shows that there exists a finite value of $\lambda$ that corresponds to the end of the regularization path of FADSG-I. However, it is difficult to further simplify it. Meanwhile, for FADSG-II, the analogous definition of $\lambda_{\max}$ corresponds to the value such that for all $\lambda\geq\lambda_{\max}$, the solution ${\mathbf{x}}^{*}(\lambda)$ is the protected subset $\mathcal{S}_{p}$. For the end of its regularization path, a finite $\lambda$ can be shown to exist, similarly to FADSG-I (see Appendix C.2). In practice, it is computationally easier to find a suitable $\lambda_{\max}$ (which we know exists owing to Propositions 6 and 10 in Appendix C) via trial and error (see Appendix G, Table 6 for values used). This enables us to bound the interval of the regularization path for a given dataset, which in turn facilitates the selection of the regularization parameter that corresponds to a target fairness level, as explained next.
@@ 4.1 Selection of Target Fairness Level
FADSG-I: For $\lambda=0$, let ${\mathbf{x}}^{*}(0)$ denote the solution of problem (2), i.e., DSG. We also denote $l_{1}:=r_{1}({\mathbf{x}}^{*}(0))\geq 0$ as the fraction of protected vertices present in the solution of DSG. This corresponds to the level of fairness (as measured by $r_{1}(\cdot)$) that is attainable by the solution of the non-fair DSG problem, and is not guaranteed to be non-zero in general. Meanwhile, we know that ${\mathbf{x}}^{*}(\lambda_{\max})$ corresponds to the densest protected subset for which $u_{1}:=r_{1}({\mathbf{x}}^{*}(\lambda_{\max}))=1$. Let $\alpha\in[l_{1},u_{1}]$ denote the desired fairness level; i.e., we wish to select $\lambda$ such that the solution of (2), ${\mathbf{x}}^{*}(\lambda)$, is the densest subgraph that satisfies $r_{1}({\mathbf{x}}^{*}(\lambda))=\alpha$. To this end, we will exploit the fact that the function $r_{1}({\mathbf{x}}^{*}(\lambda))$ is non-decreasing in $\lambda$. Consequently, if we define the function
@@
A subtle point is that for the bisection framework to provably terminate, the functions $\psi_{1}(\lambda)$ and $\psi_{2}(\lambda)$ should be continuous for every target fairness level, and for every graph. Although we cannot show this at present, we observed in our experiments that depending on the nature of the underlying Pareto-optimal frontier, some target fairness levels may be more difficult to attain than others. However, in general, the overall procedure works very well on real datasets.
@@ 5 Quantifying the Price of Fairness
FADSG-I and II trade off subgraph density in order to promote the inclusion of protected vertices. Consequently, the density of the fair subgraphs obtained via (2) and (4) is no larger than the density of their non-fair counterpart DSG; i.e., a loss in subgraph density is the price paid for ensuring fairness. The loss incurred in attaining a target fairness level $\phi$, either $\alpha$ for FADSG-I or $\delta$ for FADSG-II, can be quantified using the price of fairness.
For a given graph $\mathcal{G}$, the price of fairness (PoF) for a fixed value of $\phi$ is
$\textsc{PoF}(\mathcal{G},\phi):=\frac{\rho_{\mathcal{G}}^{*}-\rho_{\mathcal{G}}^{*}(\phi)}{\rho_{\mathcal{G}}^{*}}:=1-\frac{\rho_{\mathcal{G}}^{*}(\phi)}{\rho_{\mathcal{G}}^{*}},$ (7)
where $\rho_{\mathcal{G}}^{*}(\phi)$ is the density of the densest subgraph in $\mathcal{G}$ attaining a target fairness level $\phi$, and $\rho_{\mathcal{G}}^{*}$ is the density of the solution of DSG. Simply stated, $\textsc{PoF}(\mathcal{G},\phi)$ is the ratio of the loss in subgraph density suffered in meeting the fairness requirement $\phi$ and the maximum density $\rho_{\mathcal{G}}^{*}$ in the absence of fairness considerations. Let $\phi_{0}$ denote the fairness level of the solution of DSG (obtained with explicitly imposing fairness), then $\rho_{\mathcal{G}}^{*}(\phi_{0})=\rho_{\mathcal{G}}^{*}$, and thus $\textsc{PoF}(\mathcal{G},\phi_{0})=0$. As $\phi$ diverges from $\phi_{0}$, the resulting density $\rho_{\mathcal{G}}^{*}(\phi)$ decreases, and consequently $\textsc{PoF}(\mathcal{G},\phi)$ approaches $1$. For a desired value of $\phi$, the closer $\textsc{PoF}(\mathcal{G},\phi)$ is to $1$, the greater is the price paid (in terms of density) for extracting the desired fair dense subgraph.
@@
Similar definitions have been considered for analyzing fairness in influence maximization (Tsang et al., 2019) and graph covering (Rahmattalabi et al., 2019). To the best of our knowledge, however, we are the first to study such a notion in the context of detecting fair dense subgraphs. Analyzing the price of fairness is inherently complex due to multiple factors like the type of graph considered, the protected subgraph, and the desired level of fairness.
Figure 3: The “lollipop” graph for $n=16$. The unprotected set (blue) forms a clique on 4 vertices whereas the protected set (red) forms a path graph on 12 vertices.
@@
The “Lollipop” Graph: We base our study on a difficult example which exemplifies the underlying tension between density and fairness. Consider the “lollipop” graph (denoted as $\mathcal{L}_{n}$). In this construction, the unprotected vertices constitute a clique of size $\sqrt{n}$ ($\mathcal{C}_{\sqrt{n}}$), while the protected vertices form a path graph on $n-\sqrt{n}$ vertices ($\mathcal{P}_{n-\sqrt{n}})$, with one end connected to $\mathcal{C}_{\sqrt{n}}$ (so that the overall graph is connected). We point out that the “lollipop” graph mimics the properties of real-world graphs in the following manner: (i) the graph $\mathcal{L}_{n}$ has $O(n)$ edges, and is hence sparse, (ii) the majority of vertices ($n-\sqrt{n}$ to be precise) have low degree ($\leq 2$), while the remaining vertices have high degree (at least $\sqrt{n}-1$). Intuitively, we expect the solution of DSG for the “lollipop” graph to be the clique $\mathcal{C}_{\sqrt{n}}$, as it corresponds to a region of high density. Since this solution does not contain any protected vertices, from a fairness perspective, it is the worst possible solution. In addition, as the protected subset $\mathcal{P}_{n-\sqrt{n}}$ is minimally connected, and thus has low density, we expect that improving the representation of protected vertices will cost a large decrease in density, reflected by a large price of fairness.
@@
FADSG-I: Let us analyse how the PoF is affected by altering the balance between the majority and minority vertices in the desired solution. Consider the case where the desired proportion of unprotected and protected vertices in the extracted subgraph is $1:\gamma$, with $\gamma\in\{1,2,\cdots,n-\sqrt{n}\}$, and the desired balance is $\alpha_{\gamma}:=\gamma/(1+\gamma)$. Note that the case of $\gamma=1$ corresponds to the perfect balance scenario considered in Anagnostopoulos et al. (2020; 2024). Then, the PoF can be expressed as follows (see Appendix D.1 for proof).
@@
FADSG-II: Regarding FADSG-II, we can show that for the lollipop graph, there exists a one-to-one correspondence between the target fairness parameters $\delta$ of FADSG-II and $\alpha$ of FADSG-I such that solutions of the two problems coincide (refer to Appendix D.2 for proof). Hence the analysis remains the same as that of FADSG-I.
@@ 6 Algorithmic Approach
An alternative is to consider the iterative convex-optimization algorithms developed in Danisch et al. (2017); Harb et al. (2022); Nguyen & Ene (2024). These methods are centered around solving the dual LP relaxation of DSG (Charikar, 2000). However, modifying and implementing these algorithms (which are tailored for DSG) for our problems poses a greater challenge relative to Super-Greedy++. FADSG-I and II have to be reformulated as a constrained convex optimization problem via an extension of the dual linear programming (LP) relaxation of DSG proposed in Charikar (2000). Moreover, since the algorithms of Danisch et al. (2017); Harb et al. (2022); Nguyen & Ene (2024) are tailored to exploit the structure of the dual LP relaxation based on DSG, it is not clear to what extent they would need to be modified in order to be applicable for our problems, and whether the requisite changes would still result in low-complexity updates. Another practical drawback of these continuous methods (even when applied to plain DSG) is that a subroutine known as fractional peeling (Harb et al., 2022) has to be applied on each of the generated iterates if one desires to extract an estimate of the densest supermodular subset (which corresponds to the densest subgraph for DSG) at each iteration. However, performing a single round of fractional peeling incurs a complexity of $O(m+n\log n)$, which has the same complexity order as a single iteration of Super-Greedy++. Consequently, determining the point at which the estimate of the DSS has converged from the continuous iterates can prove to be computationally expensive, unless the fractional peeling subroutine is sparingly used. Even then, this can come at the cost of executing the convex optimization algorithms for a larger number of iterations than required for the sequence of DSS estimates to converge. We point out that this is a non-issue in Super-Greedy++, as its combinatorial nature makes it easy to keep track of the best estimate of the DSS at each iteration. Owing to these reasons, we elect to choose Super-Greedy++ as the algorithmic primitive for FADSG-I and II.
@@
Note that the objective function of FADSG-I can be expressed in the form of the DSS problem, with a non-negative, monotone supermodular numerator. Hence, the Super-Greedy++ algorithm can be applied directly to problem (2) to obtain a high-quality approximate solution. For FADSG-II, its objective function can be expressed as well as the ratio $F(\mathcal{S})/|\mathcal{S}|$, where the numerator is supermodular, making problem (4) a special case of DSS. However, since the numerator may not always be non-negative for every $\lambda>0$, Super-Greedy++ does not guarantee an $(1-\epsilon)$-approximation in all cases (see Huang et al. (2024) for a counter-example). Nonetheless, Super-Greedy++ can still be applied on problem (4) and our experiments indicate that it can serve as an effective heuristic. For details regarding implementation, refer to Appendix E.
It is worth pointing out that although we employ approximation algorithms for our problem formulations, there is a difference compared to that of the prior art Anagnostopoulos et al. (2024; 2020); Miyauchi et al. (2023). More specifically, we employ approximation as a means of improving scalability, as opposed to dealing with intractability.
@@ 7 Experimental Evaluation
Experimental Setup: We consider various real-world datasets, with Table 1 summarizing their statistics (for more information, see Appendix F). The implementation of our approach is publicly available on Github [^1] https://github.com/ekariotakis/fadsg_tmlr2025.
Table 1: Summary of dataset statistics: number of vertices ($n$), number of edges ($m$), and size of protected subset ($n_{p}$). The datasets contain multiple graphs and the statistics represent mean values with standard deviations. The full list of datasets can be found in Appendix F.
@@
  • 2-DFSG (Anagnostopoulos et al., 2020; 2024). The algorithm starts from the solution of DSG and subsequently introduces vertices from the underrepresented group in an arbitrary fashion, until perfect balance is reached.
  • PS, FPS (Anagnostopoulos et al., 2020; 2024). Paired Sweep (PS) and Fair Paired Sweep (FPS) are spectral algorithms. They categorize the elements based on their group and sort them in a manner reminiscent of spectral clustering.
@@ 7.1 Evaluation of FADSG-I
Comparison with prior art: In order to compare FADSG-I with the baselines, we set the target fairness level $\alpha=0.5$. Therefore, our formulation aims to extract a dense subgraph with perfect balance. We point out that we exclude datasets for which the protected vertices in the DSG solution are more than the unprotected, as they already meet and exceed the target $\alpha=0.5$. Figures 2 and 4 illustrate a comparative analysis with prior art. The top sections showcase the price of fairness (PoF) for each algorithm across various datasets, while the bottom sections depict the fraction of protected vertices in the solutions. As the results indicate, our proposed formulation consistently yields the lowest PoF values across a broad spectrum of scenarios. [^2] Note that, in all experiments, in order to compute $\rho^{*}_{\mathcal{G}}$, we employ the exact max flow algorithm (Goldberg, 1984).
It can be observed from Figures 2 and 4 that for certain datasets, FADSG-I does not find a perfectly balanced subgraph (see Figure 4, Amazon hpc & op). To obtain a better understanding, we plot the regularization path of $r_{1}(\mathcal{S})$ for these datasets in Figure 5. A noteworthy observation is the emergence of “fairness plateaus” where the fairness level remains constant across intervals of $\lambda$. Since FADSG-I is tractable, these figures correspond to the actual Pareto frontier which characterizes the trade-off between fairness and density. This phenomenon suggests that the graph’s inherent structure leads to certain fairness levels that arise naturally. For instances where the data’s underlying structure does not inherently support a perfect division (i.e., Amazon hpc & op), FADSG-I does not enforce it. Figure 6 depicts the outcomes in datasets where there exists an extreme imbalance in the composition of the protected and unprotected subgroups, i.e., we have $n_{p}\ll n$ (see Table 4, Appendix F). In such cases, DDSP consistently fails to produce balanced/fair solutions. In fact, we observe an extreme gap to the desired fairness level of $0.5$, as well as a PoF that is always worse than that of FADSG-I. The other prior methods always achieve a perfect balance, but with a significantly higher PoF compared to our algorithm. This is an important feature of FADSG-I, as it can provide a near optimal solution, when the nature of the graph allows it, while almost always paying the least in terms of PoF.